# Learning Optimal Control with Stochastic Models of Hamiltonian Dynamics

### v2

@@ 4.2 Second phase of variational training
[…]
@@
[…]
@@
The training process is similar to that of a variational auto-encoder Kingma and Welling (2019) and we wish to optimize the loss function:
@@
[…]
@@ 5.1 Models
    Random Hamiltonian model outputs a random action, which nonetheless follows Hamiltonian dynamics bias rather than totally random.
@@
    NeuralPMP-phase 1 is the model that is trained with only the first phase $Forward(T)$ with deterministic reduced Hamiltonian and adjoint neural networks training.
@@
    NeuralPMP, our main framework, that is described in details in the previous section section 4 with 2 phases $Forward(T)$ and $Backward(T)$
@@ 5.2 Environments:
The smaller the evaluation cost, the better the agents get. All 3 problems are continuous-versions of environments from the OpenAI Gym suite (Brockman et al. (2016)).
@@ 5.3 Evaluation
We evaluate 3 models by running each model on $2000$ initial starting points to collect $2000$ trajectories. For each of such trajectories, we calculate the evaluation cost at each time steps between 00 and terminal time $T=1$ with step size $0.05$ (21 time steps in total). For each model and at each time step, we then calculate the mean evaluation cost over $2000$ trajectories. We also use bootstrap method to calculate the adjusted variance of the evaluation costs on those $2000$ trajectories by performing 2 steps:
@@
The plots fig. 4 show how evaluation costs (mean and adjusted variance on 2000 trajectories) vary across different time steps for each model on each problem. The plots show that both NeuralPMP models outperform the random Hamiltonian dynamics model. The NeuralPMP model performs slightly better than the version with only phase 1 (forward) training. Even in the pendulum case, even though the cost is higher at terminal time for NeuralPMP, such cost still drops faster in most time steps than the one from one phase-model. This is due to the stochasticity added via the forward-backward Hamiltonian flows variational auto-encoder structure. This structure overall improves exploration and generalizability while learning the Hamiltonian network accurately. We also report in the table 1 the average evaluation costs at the terminal time $T=1$ of 3 models on 3 problems. Details about hyper-parameters, neural network architectures, and the continuous-time design of each problem are given the the appendix.
@@ 6 Conclusion
In this paper, we build learning frameworks for optimal control problems by actively using the Pontryagin maximum principle for training and learning rather than for mathematical purposes. We first present an optimal planning discrete-time algorithm. We then present our main two-phases learning framework called NeuralPMP that allows learning a reduced Hamiltonian dynamics and at the same time derives the optimal control. In the future, we plan to develop model-free versions of our framework to eliminate the need to access black-box dynamics or black-box running/terminal costs. We also would like to extend the framework to larger-scale problems with higher-dimensional states and actions.
Author acknowledgement: This research was supported in part by a grant from the NIH DK129979, in part from the Peter O’Donnell Foundation, the Michael J. Fox Foundation, Jim Holland-Backcountry Foundation, and in part from a grant from the Army Research Office accomplished under Cooperative Agreement Number W911NF-19-2-0333.
Author Contributions: NeuralPMP is developed by Minh Nguyen and improved by Chandrajit Bajaj. The implementation is done by Minh Nguyen and is available at NeuralPMP project.
@@ Appendix B Details of costs and dynamics for control problems
For the classical control tasks: Cart pole, Mountain car, and Pendulum, we choose the dynamics similar to the one given in the OpenAI Gym suite (Brockman et al. (2016)). The states for these two problems are:

### v3

@@ 4.2 Second phase of variational training
More concretely, after the first training procedure $\textit{Forward}(T)$ is finished, a reverse Hamiltonian flow given by another parameterized reduced Hamiltonian $h^{decoder}_{\lambda}=h^{decoder}_{\lambda}(q,p)$ will be learned. Such learning process include 3 steps:
@@
    We first variationally encode $y:=(q_{T},p_{T})$ from the $\textit{Forward}(T)$ phase into a latent variable $z\in\mathbb{R}^{d}$, for a latent dimension $d\in\mathbb{N}$ via an encoder $E_{\lambda_{1}}$ so that $z=E_{\lambda_{1}}(y)=E_{\lambda_{1}}(q_{T},p_{T})$.
@@
The training process is similar to that of a variational auto-encoder [13] and we optimize the following loss function:
@@
The first two terms are reconstruction errors with $y=(q_{T},p_{T})$, $\hat{y}=(\widehat{q_{T}},\widehat{p_{T}})$, $x=(q_{0},p_{0})$, $\hat{x}=(\widehat{q_{0}},\widehat{p_{0}})$. The last term is the usual KL-divergence with $P$ and $Q$ being the prior and posterior distributions of the latent variable $z$. The stochasticity now comes from the sampling of $z$ during the training process. $\beta_{2}$ is a hyperparameter that is again found by grid search with early termination. Note that the reverse flow leading to $(q_{0},p_{0})$ is based on the backward ordinary differential equation with the dynamics function $(-h^{decoder}_{\lambda})$. We call this additional training phase $\textit{Backward}(T)$ for the terminal time $T$. The final NeuralPMP training procedure includes two phases: Forward(T) and then Backward(T) trained sequentially with the same terminal time $T$.
@@ 5.1 Models
    Random Hamiltonian: This model outputs a random action, which nonetheless follows Hamiltonian dynamics bias rather than totally random.
@@
    NeuralPMP-phase 1: This model is trained with only the first phase $Forward(T)$ with deterministic reduced Hamiltonian and adjoint neural networks training.
@@
    NeuralPMP: This is our main framework described in details in the previous section 4 with 2 phases $Forward(T)$ and $Backward(T)$.
@@ 5.2 Environments:
The smaller the evaluation cost, the better the agents perform. All 3 problems are continuous versions of the environments from the OpenAI Gym suite [5].
@@ 5.3 Evaluation
We evaluate 3 models by running each model on $2000$ initial starting points to collect $2000$ trajectories. For each of such trajectories, we calculate the evaluation cost at each time step between 00 and terminal time $T=1$ with step size $0.05$ (21 time steps in total). For each model and at each time step, we calculate the mean evaluation cost over $2000$ trajectories. We also use bootstrap method to calculate the adjusted variance of the evaluation costs on those $2000$ trajectories by performing 2 steps:
@@
The plots (see fig. 4) show how evaluation costs (mean and adjusted variance on 2000 trajectories) vary across different time steps for each model on each problem. The plots show that both NeuralPMP models outperform the random Hamiltonian dynamics model. NeuralPMP model performs slightly better than the version with only phase 1 (forward) training. Even in the pendulum case, while the cost is higher at terminal time for NeuralPMP, such cost still drops faster in most time steps than the one from one phase-model. This is due to the stochasticity added via the forward-backward Hamiltonian flows variational auto-encoder structure. This structure overall improves exploration and generalizability while learning the Hamiltonian network accurately. We also report in table 1 the average evaluation costs at the terminal time $T=1$ of 3 models on 3 problems. Details about hyper-parameters, neural network architectures, and the continuous-time design of each problem are provided in the appendix.
@@ 6 Conclusion
In this paper, we build a learning framework for optimal control problems by actively using the Pontryagin maximum principle for training and learning and beyond just invocation of a mathematical principle. We first present an optimal planning discrete-time algorithm. We then introduce our main two-phase learning framework called NeuralPMP. This framework allows learning a reduced Hamiltonian dynamics and at the same time derives the optimal control or actions. In the future, we plan to develop model-free versions of our framework to eliminate the need to access black-box dynamics or black-box running/terminal costs. We also would like to extend the framework to larger-scale problems with higher-dimensional states and actions.
Author acknowledgement: This research work was supported in part by a grant from the NIH DK129979, in part from the Peter O’Donnell Foundation, the Michael J. Fox Foundation, Jim Holland-Backcountry Foundation, and in part from a grant from the Army Research Office accomplished under Cooperative Agreement Number W911NF-19-2-0333.
Author Contributions: This research work is developed by Minh Nguyen and improved by Chandrajit Bajaj. Alphabetical order is currently used to order author names. The implementation is done by Minh Nguyen and is available at NeuralPMP project.
@@ Appendix B Details of costs and dynamics for control problems
For the classical control tasks: Cart pole, Mountain car, and Pendulum, we choose the dynamics similar to the one given in the OpenAI Gym suite [5]. The states for these two problems are: